# Geometry of the moduli space of Hermitian-Einstein connections on manifolds with a dilaton

### v1

@@ 4 Symmetries of moduli spaces and the dilaton field
$\mathcal{L}_{\alpha_{{}_{X}}}(\alpha_{1},\alpha_{2})_{{\mathscr{M}}^{*}_{\mathrm{HE}}}=\int_{M^{2n}}d^{2n}x\sqrt{g}\,\,e^{-2\Phi}\,\,g^{-1}\mathcal{L}_{X}\langle a_{1}^{h}\,,\,a_{2}^{h}\rangle_{\mathfrak{g}}=0\leavevmode\nobreak\ ,$ (4.1)
@@
$\mathcal{L}_{\alpha_{{}_{X}}}\Omega(\alpha_{1},\alpha_{2})=\frac{1}{(n-1)!}\int_{M^{2n}}\,\,e^{-2\Phi}\,\,\omega^{n-1}\wedge\mathcal{L}_{X}\langle a_{1}^{h}\wedge a^{h}_{2}\rangle_{\mathfrak{g}}=0\leavevmode\nobreak\ ,$ (4.2)
@@
It remains to investigate the special case that $X$ is $\widehat{\nabla}$ parallel and holomorphic. The former conditions implies that $X$ is Killing and $dX^{\flat}=\iota_{X}H$. Conversely, if $X$ is Killing and $dX^{\flat}=\iota_{X}H$, then $X$ is $\widehat{\nabla}$-covariantly constant. This will be used in the theorem below to demonstrate that $\alpha_{X}$ is $\widehat{\mathcal{D}}$-covariantly constant on $\mathscr{M}^{*}_{\mathrm{HE}}$. Assumming that $(M^{2n},g,\omega)$ is a compact KT manifold with dilaton $\Phi$ such that the condition (2.20), $D^{i}\big{(}e^{-2\Phi}K^{\flat}_{i}\big{)}=0$ with $K^{\flat}=\theta+2d\Phi$, is satisfied, one can establish the following result.
@@
To begin, it is a consequence of lemma 3.2 and in particular (3.6) that $a_{X}^{h}=\iota_{X}F$ as ^$\widehat{\nabla}X=0$. To continue, we have to compute $\iota_{\alpha_{X}}\mathcal{H}$. Indeed, a computation similar to that presented in [27] reveals that
@@
$\alpha_{{}_{X}}^{\flat}(\alpha)\equiv\mathcal{F}_{X}(a^{h})\equiv\int_{M^{2n}}\sqrt{g}\,\,e^{-2\Phi}\,\,g^{-1}\langle\iota_{X}F,a^{h}\rangle_{\mathfrak{g}}\leavevmode\nobreak\ .$ (4.5)
@@
$d\mathcal{F}_{X}(a^{h}_{2},a^{h}_{3})=\int_{M^{2n}}d^{2n}x\sqrt{g}\,\,e^{-2\Phi}\,\,g^{-1}\big{(}\langle\iota_{X}d_{A}a_{2}^{h},a_{3}^{h}\rangle_{\mathfrak{g}}-\langle\iota_{X}d_{A}a_{3}^{h},a_{2}^{h}\rangle_{\mathfrak{g}}-\langle\iota_{X}F,d_{A}\Theta(a_{2}^{h},a_{3}^{h})\rangle_{\mathfrak{g}}\big{)}\leavevmode\nobreak\ .$ (4.6)
@@
$d\mathcal{F}_{X}(a^{h}_{2},a^{h}_{3})$ $=\int_{M^{2n}}d^{2n}x\sqrt{g}\,\,e^{-2\Phi}\,\,\big{(}(IX\wedge IK)^{ij}\langle a_{2i}^{h},a_{3j}^{h}\rangle_{\mathfrak{g}}-g^{-1}\langle\iota_{X}F,d_{A}\Theta(a_{2}^{h},a_{3}^{h})\rangle_{\mathfrak{g}}\big{)}\leavevmode\nobreak\ .$ (4.9)
@@
$X^{\flat}\wedge K^{\flat}=\iota_{I}X^{\flat}\wedge\iota_{I}K^{\flat}\leavevmode\nobreak\ ,$ (4.10)
@@ 5 Holomorphic and Killing vector fields on geometries with a dilaton field
$D^{i}\big{(}e^{-2\Phi}K^{\flat}_{i}\big{)}=0\leavevmode\nobreak\ ,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ K^{\flat}=\theta+2d\Phi\leavevmode\nobreak\ .$ (5.1)
@@
$\widehat{R}_{ij}+2\widehat{\nabla}_{i}\partial_{j}\Phi-{1\over 4}P_{ij}=0\leavevmode\nobreak\ ,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak$ (5.2)
for $P_{ij}=P_{ji}$ that will be specified below. As we have already explained, these geometries include all those that have found applications in either string theory or in the theory of geometric flows. The condition on the geometry (5.2) is one of the assumptions motivated by the applications mentioned above. We shall not assume that $M^{d}$ is compact unless it is explicitly stated.
@@ 5.1 Symmetries of KT geometry with dilaton
$\mathcal{L}_{K}g^{2,0}=0\leavevmode\nobreak\ ,$ (5.3)
$\mathcal{L}_{K}\omega^{2,0}=0\leavevmode\nobreak\ ,$ (5.4)
$\mathcal{L}_{K}\omega=-2\rho^{1,1}\leavevmode\nobreak\ ,$ (5.5)
$\mathcal{L}_{Y}g=\mathcal{L}_{Y}\omega=0,\leavevmode\nobreak\ \mathrm{where}\leavevmode\nobreak\ \leavevmode\nobreak\ Y=-IK\leavevmode\nobreak\ ,$ (5.6)
@@
In particular, $K$ is holomorphic, but not necessarily Killing, vector field and $Y=-IK$ is both holomorphic and Killing vector field.
@@
$\widehat{R}_{ikmn}+\mathrm{cyclic\leavevmode\nobreak\ in}\leavevmode\nobreak\ \leavevmode\nobreak\ (k,\,m,\,n)=-\widehat{\nabla}_{i}H_{kmn}+\frac{1}{2}dH_{ikmn}\leavevmode\nobreak\ .$ (5.7)
@@
$\rho_{ik}I^{k}{}_{j}+\widehat{R}_{ij}-\frac{1}{4}P_{ij}=\widehat{\nabla}_{i}\theta_{j}\Longrightarrow\rho_{ik}I^{k}{}_{j}=\widehat{\nabla}_{i}(\theta_{j}+2\partial_{j}\Phi)=\widehat{\nabla}_{i}K_{j}\leavevmode\nobreak\ ,$ (5.8)
@@
$\rho_{ij}=\frac{1}{2}\widehat{R}_{ijmn}\omega^{mn}\leavevmode\nobreak\ .$ (5.9)
@@
^$\widehat{\nabla}_{\alpha}K_{\beta}+\widehat{\nabla}_{\beta}K_{\alpha}=\nabla_{\alpha}K_{\beta}+\nabla_{\beta}K_{\alpha}=i\rho_{\alpha\beta}+i\rho_{\beta\alpha}=0$ (5.10)
as $P$ is a $(1,1)$ tensor, where we have used complex coordinates $(x^{i};i=1,\dots,2n)=(z^{\alpha},z^{\bar{\alpha}};\alpha=1,\dots,n)$ with $I=(i\delta^{\alpha}{}_{\beta},-i\delta^{\bar{\alpha}}{}_{\bar{\beta}})$. This proves the first statement.
@@
$H_{\alpha\bar{\beta}\bar{\gamma}}=-\partial_{\bar{\beta}}g_{\alpha\bar{\gamma}}+\partial_{\bar{\gamma}}g_{\alpha\bar{\beta}}\leavevmode\nobreak\ ,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \theta_{\alpha}=H_{\alpha\beta\bar{\gamma}}g^{\beta\bar{\gamma}}\leavevmode\nobreak\ .$ (5.11)
@@
$D^{i}H_{i\beta\gamma}=g^{\alpha\bar{\alpha}}D_{\alpha}H_{\bar{\alpha}\beta\gamma}+g^{\alpha\bar{\alpha}}D_{\bar{\alpha}}H_{\alpha\beta\gamma}=-\partial_{\beta}\theta_{\gamma}+\partial_{\gamma}\theta_{\beta}-\theta_{\delta}H^{\delta}{}_{\beta\gamma}\leavevmode\nobreak\ ,$ (5.12)
@@
$\mathcal{L}_{\theta^{\flat}}\omega_{\beta\gamma}=i\partial_{\beta}\theta_{\gamma}-i\partial_{\gamma}\theta_{\beta}+i\theta_{\delta}H^{\delta}{}_{\beta\gamma}\leavevmode\nobreak\ .$ (5.13)
@@
$D^{i}H_{i\beta\gamma}=i\mathcal{L}_{\theta^{\flat}}\omega_{\beta\gamma}\leavevmode\nobreak\ .$ (5.14)
@@
$D^{i}H_{i\beta\gamma}=2H^{\delta}{}_{\beta\gamma}\partial_{\delta}\Phi\leavevmode\nobreak\ .$ (5.15)
@@
Notice that in complex coordinates, $I=(i\delta^{\alpha}{}_{\beta},-i\delta^{\bar{\alpha}}{}_{\bar{\beta}})$, the first and second properties imply that $\partial_{\bar{\beta}}K^{\alpha}=0$, i.e. that $K$ is holomorphic, and so
$\mathcal{L}_{K}I=0\leavevmode\nobreak\ .$ (5.17)
@@
$\mathcal{L}_{K}\omega_{\alpha\bar{\beta}}=-i\mathcal{L}_{K}g_{\alpha\bar{\beta}}=-2\rho_{\alpha\bar{\beta}}\leavevmode\nobreak\ .$ (5.19)
Next as $K$ is holomorphic so is $Y=-IK$. It remains to show that it is Killing. This follows from (5.8) as
$\rho_{ij}=-\widehat{\nabla}_{i}(K^{\flat}_{m}I^{m}{}_{j})=-\widehat{\nabla}_{i}Y^{\flat}_{j}\leavevmode\nobreak\ .$ (5.20)
@@
Suppose that $(M^{2n},g,\omega)$ is a compact KT manifold that satisfies (5.2) and $Z$ is a holomorphic and Killing vector field. Then, the dilaton field is invariant under the action of $Z$, $Z\Phi=0$. In particular, $Y$ leaves $\Phi$ invariant, $Y\Phi=0$, for $Y=-IK$.
@@
In proposition 5.1, we have demonstrated that $Y=-IK$ is both holomorphic and Killing. Then, it follows from the argument above that $Y\Phi=0$ and so $\Phi$ is invariant under the action of $Y$. ∎
@@ 5.1.1 Symmetries of CYT and HKT geometries with dilaton
Suppose that $(M^{2n},g,\omega)$ is a CYT manifold that satisfies (5.2), then $K$ and $Y=-IK$ with $K^{\flat}=\theta+2d\Phi$ are both holomorphic and $\widehat{\nabla}$-covariantly constant. In particular, both $K$ and $Y$ are Killing vector fields.
@@
^$\widehat{\nabla}K=0\leavevmode\nobreak\ ,$ (5.21)
i.e. that $K$ is $\widehat{\nabla}$-covariantly constant. It follows from this that $K$ is Killing and $dK^{\flat}=\iota_{K}H$.
We have already demonstrated that $Y$ is holomorphic and Killing. It is also $\widehat{\nabla}$-covariantly constant as both $I$ and $K$ are $\widehat{\nabla}$-covariantly constant. ∎
@@
As $K$ and $Y=-IK$ are holomorphic and $\widehat{\nabla}$-covariantly constant, it follows that they commute $[K,Y]=0$. This has been shown in [46]. This is a consequence of the fact that the commutator of two $\widehat{\nabla}$-covariantly constant vector field can be expressed in terms of $\iota_{K}\iota_{Y}H$ and that $\iota_{K}H$ is a $(1,1)$-form as $K$ is holomorphic. Furthermore, if $M^{2n}$ is compact, then as a consequence of proposition 5.2 one has $K\Phi=Y\Phi=0$.
@@
$(\rho_{r})_{ij}=\frac{1}{2}\widehat{R}_{ikmn}I^{k}_{rj}\omega^{mn}_{r}\leavevmode\nobreak\ ,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ (\mathrm{no\leavevmode\nobreak\ summation\leavevmode\nobreak\ over\leavevmode\nobreak\ r})\leavevmode\nobreak\ ,$ (5.22)
@@
$\theta=\theta_{1}=\theta_{2}=\theta_{3}\leavevmode\nobreak\ ,$ (5.23)
@@
Suppose that $(M^{4k},g,\omega_{r})$ is an HKT manifold that satisfies (5.2), then $K$ with $K^{\flat}=\theta+2d\Phi$ is tri-holomorphic and $\widehat{\nabla}$-covariantly constant vector field on $M^{4k}$, i.e. $K$ is holomorphic with respect to all three complex structures. Moreover, $Y_{r}=-I_{r}K$ is holomorphic with respect to the $I_{r}$ complex structure and $\widehat{\nabla}$-covariantly constant. In particular, the vector fields $K$ and $Y_{r}$ are Killing.
@@
It is clear from the definition of HKT manifolds that they admit a CYT structure with respect to each one of the complex structures $I_{r}$. As the Lee form of the geometry does not depend on the choice of the complex structure that it is used to define it, it follows from the previous corollary that $K$ must be holomorphic with respect to each one of the complex structures $I_{r}$ and $\widehat{\nabla}$-covariantly constant. Thus, in particular, $K$ is tri-holomorphic.
It also follows from the previous corollary that $Y_{r}$ must be holomorphic with respect to the $I_{r}$ complex structure. It is also $\widehat{\nabla}$-covariantly constant as both $I_{r}$ and $K$ are $\widehat{\nabla}$-covariantly constant. The $\widehat{\nabla}$-covariant constancy condition implies that all $K$ and $Y_{r}$ are Killing vector fields. ∎
@@
A similar argument to that deployed in the CYT case above implies that $K$ commutes with $Y_{r}$, i.e. $[K,Y_{r}]=0$. However in general $[Y_{r},Y_{s}]\not=0$, for $r\not=s$. An example of such symmetries is the Lie algebra generated by the left- or right-invariant vector fields of HKT geometry on $S^{3}\times S^{1}$ viewed as the group manifold $SU(2)\times U(1)$. In this example, $K=\partial_{\tau}$ with $\tau$ the coordinate along $S^{1}$ and $Y_{r}=-I_{r}K$ are the left-invariant vector fields on $S^{3}$, where $I_{r}$ are the left-invariant complex structures on $S^{3}\times S^{1}$, see [23] for a detailed analysis.
@@
If in addition the HKT manifold $M^{4k}$ is compact, then as a consequence of proposition 5.2, one has that $K\Phi=Y_{r}\Phi=0$. In four dimensions, $M^{4}$ is parallelisable because $K$ and $Y_{r}$ are also $\widehat{\nabla}$-covariantly constant. Since $K$ commutes with $Y_{r}$ and the commutator is given by the components of $H$, $H$ can be written as $H=f\,Y^{\flat}_{1}\wedge Y^{\flat}_{2}\wedge Y^{\flat}_{3}$ for some function $f$. However, $H$ is invariant under the action of $K$ and $Y_{r}$, because $K$ is tri-holomorphic and Killing and $Y_{r}$ is holomorphic with respect to $I_{r}$ and Killing, which implies that $f$ is constant and $H$ is necessarily a closed 3-form. As $\Phi$ is invariant under the action of all four vectors fields $K$ and $Y_{r}$, $\Phi$ is constant. These manifolds for $H\not=0$ are identifications of $S^{3}\times S^{1}$ with a discrete group. Another way to state this result is that the unique set of hyper-complex structures on $S^{3}\times S^{1}$ [50] do not admit HKT geometries with $dH\not=0$ and non-constant dilaton $\Phi$. For a detailed description of the HKT structures on $S^{3}\times S^{1}$ see [23].

### v2

@@ 4 Symmetries of moduli spaces and the dilaton field
$\mathcal{L}_{\alpha_{{}_{X}}}(\alpha_{1},\alpha_{2})_{{\mathscr{M}}^{*}_{\mathrm{HE}}}=\int_{M^{2n}}d^{2n}x\sqrt{g}\,\,e^{-2\Phi}\,\,g^{-1}\mathcal{L}_{X}\langle a_{1}^{h}\,,\,a_{2}^{h}\rangle_{\mathfrak{g}}=0\penalty 10000\ ,$ (4.1)
@@
$\mathcal{L}_{\alpha_{{}_{X}}}\Omega(\alpha_{1},\alpha_{2})=\frac{1}{(n-1)!}\int_{M^{2n}}\,\,e^{-2\Phi}\,\,\omega^{n-1}\wedge\mathcal{L}_{X}\langle a_{1}^{h}\wedge a^{h}_{2}\rangle_{\mathfrak{g}}=0\penalty 10000\ ,$ (4.2)
@@
It remains to investigate the special case that $X$ is $^\widehat{\nabla}$-covariantly constant and holomorphic. The former conditions implies that $X$ is Killing and $dX^{\flat}=\iota_{X}H$. Conversely, if $X$ is Killing and $dX^{\flat}=\iota_{X}H$, then $X$ is $^\widehat{\nabla}$-covariantly constant. This will be used in the theorem below to demonstrate that $\alpha_{X}$ is $\widehat{\mathcal{D}}-$covariantly constant on $\mathscr{M}^{*}_{\mathrm{HE}}$. Assumming that $(M^{2n},g,\omega)$ is a compact KT manifold with dilaton $\Phi$ such that the condition (2.20), $D^{i}\big(e^{-2\Phi}K^{\flat}_{i}\big)=0$ with $K^{\flat}=\theta+2d\Phi$, is satisfied, one can establish the following result.
@@
To begin, it is a consequence of Lemma 3.2 and in particular (3.6) that $a_{X}^{h}=\iota_{X}F$ as $\widehat{\nabla}X=0$. To continue, we have to compute $\iota_{\alpha_{X}}\mathcal{H}$. Indeed, a computation similar to that presented in [27] reveals that
@@
$\alpha_{{}_{X}}^{\flat}(\alpha)\equiv\mathcal{F}_{X}(a^{h})\equiv\int_{M^{2n}}\sqrt{g}\,\,e^{-2\Phi}\,\,g^{-1}\langle\iota_{X}F,a^{h}\rangle_{\mathfrak{g}}\penalty 10000\ .$ (4.5)
@@
$d\mathcal{F}_{X}(a^{h}_{2},a^{h}_{3})=\int_{M^{2n}}d^{2n}x\sqrt{g}\,\,e^{-2\Phi}\,\,g^{-1}\big(\langle\iota_{X}d_{A}a_{2}^{h},a_{3}^{h}\rangle_{\mathfrak{g}}-\langle\iota_{X}d_{A}a_{3}^{h},a_{2}^{h}\rangle_{\mathfrak{g}}-\langle\iota_{X}F,d_{A}\Theta(a_{2}^{h},a_{3}^{h})\rangle_{\mathfrak{g}}\big)\penalty 10000\ .$ (4.6)
@@
$d\mathcal{F}_{X}(a^{h}_{2},a^{h}_{3})$ $=\int_{M^{2n}}d^{2n}x\sqrt{g}\,\,e^{-2\Phi}\,\,\big((IX\wedge IK)^{ij}\langle a_{2i}^{h},a_{3j}^{h}\rangle_{\mathfrak{g}}-g^{-1}\langle\iota_{X}F,d_{A}\Theta(a_{2}^{h},a_{3}^{h})\rangle_{\mathfrak{g}}\big)\penalty 10000\ .$ (4.9)
@@
$X^{\flat}\wedge K^{\flat}=\iota_{I}X^{\flat}\wedge\iota_{I}K^{\flat}\penalty 10000\ ,$ (4.10)
@@ 5 Holomorphic and Killing vector fields on geometries with a dilaton field
$D^{i}\big(e^{-2\Phi}K^{\flat}_{i}\big)=0\penalty 10000\ ,\penalty 10000\ \penalty 10000\ \penalty 10000\ K^{\flat}=\theta+2d\Phi\penalty 10000\ .$ (5.1)
@@
$\widehat{R}_{ij}+2\widehat{\nabla}_{i}\partial_{j}\Phi-{1\over 4}P_{ij}=0\penalty 10000\ ,\penalty 10000\ \penalty 10000\ \penalty 10000\$(5.2)
for $P_{ij}=P_{ji}$ that will be specified below. As we have already explained, these geometries include all those that have found applications in either string theory or in the theory of geometric (generalised) flows. The condition on the geometry (5.2) is one of the assumptions motivated by the applications mentioned above. We shall not assume that $M^{d}$ is compact unless it is explicitly stated.
@@ 5.1 Symmetries of KT geometry with dilaton
$\mathcal{L}_{K}g^{2,0}=0\penalty 10000\ ,$ (5.3)
$\mathcal{L}_{K}\omega^{2,0}=0\penalty 10000\ ,$ (5.4)
$\mathcal{L}_{K}\omega=-2\rho^{1,1}\penalty 10000\ ,$ (5.5)
$\mathcal{L}_{Y}g=\mathcal{L}_{Y}\omega=0,\penalty 10000\ \mathrm{where}\penalty 10000\ \penalty 10000\ Y=-IK\penalty 10000\ ,$ (5.6)
@@
In particular, $K$ is holomorphic, but not necessarily Killing, vector field and Y=IKY=-IK is both holomorphic and Killing vector field.
@@
$\widehat{R}_{ikmn}+\mathrm{cyclic\penalty 10000\ in}\penalty 10000\ \penalty 10000\ (k,\,m,\,n)=-\widehat{\nabla}_{i}H_{kmn}+\frac{1}{2}dH_{ikmn}\penalty 10000\ .$ (5.7)
@@
$\rho_{ik}I^{k}{}_{j}+\widehat{R}_{ij}-\frac{1}{4}P_{ij}=\widehat{\nabla}_{i}\theta_{j}\Longrightarrow\rho_{ik}I^{k}{}_{j}=\widehat{\nabla}_{i}(\theta_{j}+2\partial_{j}\Phi)=\widehat{\nabla}_{i}K_{j}\penalty 10000\ ,$ (5.8)
@@
$\rho_{ij}=\frac{1}{2}\widehat{R}_{ijmn}\omega^{mn}\penalty 10000\ .$ (5.9)
@@
$\widehat{\nabla}_{\alpha}K_{\beta}+\widehat{\nabla}_{\beta}K_{\alpha}=\nabla_{\alpha}K_{\beta}+\nabla_{\beta}K_{\alpha}=i\rho_{\alpha\beta}+i\rho_{\beta\alpha}=0$ (5.10)
[…]
@@
$H_{\alpha\bar{\beta}\bar{\gamma}}=-\partial_{\bar{\beta}}g_{\alpha\bar{\gamma}}+\partial_{\bar{\gamma}}g_{\alpha\bar{\beta}}\penalty 10000\ ,\penalty 10000\ \penalty 10000\ \penalty 10000\ \theta_{\alpha}=H_{\alpha\beta\bar{\gamma}}g^{\beta\bar{\gamma}}\penalty 10000\ .$ (5.11)
@@
$D^{i}H_{i\beta\gamma}=g^{\alpha\bar{\alpha}}D_{\alpha}H_{\bar{\alpha}\beta\gamma}+g^{\alpha\bar{\alpha}}D_{\bar{\alpha}}H_{\alpha\beta\gamma}=-\partial_{\beta}\theta_{\gamma}+\partial_{\gamma}\theta_{\beta}-\theta_{\delta}H^{\delta}{}_{\beta\gamma}\penalty 10000\$, (5.12)
@@
$\mathcal{L}_{\theta^{\flat}}\omega_{\beta\gamma}=i\partial_{\beta}\theta_{\gamma}-i\partial_{\gamma}\theta_{\beta}+i\theta_{\delta}H^{\delta}{}_{\beta\gamma}\penalty 10000\$. (5.13)
@@
$D^{i}H_{i\beta\gamma}=i\mathcal{L}_{\theta^{\flat}}\omega_{\beta\gamma}\penalty 10000\ .$ (5.14)
@@
$D^{i}H_{i\beta\gamma}=2H^{\delta}{}_{\beta\gamma}\partial_{\delta}\Phi\penalty 10000\ .$ (5.15)
@@
[…]
$\mathcal{L}_{K}I=0\penalty 10000\ .$ (5.17)
@@
$\mathcal{L}_{K}\omega_{\alpha\bar{\beta}}=-i\mathcal{L}_{K}g_{\alpha\bar{\beta}}=-2\rho_{\alpha\bar{\beta}}\penalty 10000\ .$ (5.19)
Next as $K$ is holomorphic so is Y=IKY=-IK. It remains to show that it is Killing. This follows from (5.8) as
$\rho_{ij}=-\widehat{\nabla}_{i}(K^{\flat}_{m}I^{m}{}_{j})=-\widehat{\nabla}_{i}Y^{\flat}_{j}\penalty 10000\ .$ (5.20)
@@
Suppose that $(M^{2n},g,\omega)$ is a compact KT manifold that satisfies (5.2) and $Z$ is a holomorphic and Killing vector field. Then, the dilaton field is invariant under the action of $Z$, $Z\Phi=0$. In particular, $Y$ leaves $\Phi$ invariant, $Y\Phi=0$, for Y=IKY=-IK.
@@
In Proposition 5.1, we have demonstrated that Y=IKY=-IK is both holomorphic and Killing. Then, it follows from the argument above that $Y\Phi=0$ and so $\Phi$ is invariant under the action of $Y$. ∎
@@ 5.1.1 Symmetries of CYT and HKT geometries with dilaton
Suppose that $(M^{2n},g,\omega)$ is a CYT manifold that satisfies (5.2), then $K$ and Y=IKY=-IK with $K^{\flat}=\theta+2d\Phi$ are both holomorphic and $^\widehat{\nabla}$-covariantly constant. In particular, both $K$ and $Y$ are Killing vector fields.
@@
$\widehat{\nabla}K=0\penalty 10000\ ,$ (5.21)
i.e. that $K$ is $^\widehat{\nabla}$-covariantly constant. It follows from this that $K$ is Killing and $dK^{\flat}=\iota_{K}H$.
We have already demonstrated that $Y$ is holomorphic and Killing. It is also $^\widehat{\nabla}$-covariantly constant as both $I$ and $K$ are $^\widehat{\nabla}$-covariantly constant. ∎
@@
As $K$ and Y=IKY=-IK are holomorphic and $^\widehat{\nabla}$-covariantly constant, it easily follows that they commute $[K,Y]=0$, see e.g. [47]. This is a consequence of the fact that the commutator of two $^\widehat{\nabla}$-covariantly constant vector field can be expressed in terms of $\iota_{K}\iota_{Y}H$ and that $\iota_{K}H$ is a $(1,1)$-form as $K$ is holomorphic. Furthermore, if $M^{2n}$ is compact, then as a consequence of Proposition 5.2 one has $K\Phi=Y\Phi=0$.
@@
$(\rho_{r})_{ij}=\frac{1}{2}\widehat{R}_{ikmn}I^{k}_{rj}\omega^{mn}_{r}\penalty 10000\ ,\penalty 10000\ \penalty 10000\ \penalty 10000\ (\mathrm{no\penalty 10000\ summation\penalty 10000\ over\penalty 10000\ r})\penalty 10000\ ,$ (5.22)
@@
$\theta=\theta_{1}=\theta_{2}=\theta_{3}\penalty 10000\ ,$ (5.23)
@@
Suppose that $(M^{4k},g,\omega_{r})$ is an HKT manifold that satisfies (5.2), then $K$ with $K^{\flat}=\theta+2d\Phi$ is tri-holomorphic and $^\widehat{\nabla}$-covariantly constant vector field on $M^{4k}$, i.e. $K$ is holomorphic with respect to all three complex structures. Moreover, $Y_{r}=-I_{r}K$ is holomorphic with respect to the $I_{r}$ complex structure and $^\widehat{\nabla}$-covariantly constant. In particular, the vector fields $K$ and $Y_{r}$ are Killing.
@@
It is clear from the definition of HKT manifolds that they admit a CYT structure with respect to each one of the complex structures $I_{r}$. As the Lee form of the geometry does not depend on the choice of the complex structure that it is used to define it, it follows from the previous corollary that $K$ must be holomorphic with respect to each one of the complex structures $I_{r}$ and $^\widehat{\nabla}$-covariantly constant. Thus, in particular, $K$ is tri-holomorphic.
It also follows from the previous corollary that $Y_{r}$ must be holomorphic with respect to the $I_{r}$ complex structure. It is also $^\widehat{\nabla}$-covariantly constant as both $I_{r}$ and $K$ are $^\widehat{\nabla}$-covariantly constant. The $^\widehat{\nabla}$-covariant constancy condition implies that all $K$ and $Y_{r}$ are Killing vector fields. ∎
@@
A similar argument to that deployed in the CYT case above implies that $K$ commutes with $Y_{r}$, i.e. $K,Y_{r}]=0$. However in general $Y_{r},Y_{s}]\not=0$, for $r\not=s$. An example of such symmetries is the Lie algebra generated by the left- or right-invariant vector fields of HKT geometry on $S^{3}\times S^{1}$ viewed as the group manifold $SU(2)\times U(1)$. In this example, $K=\partial_{\tau}$ with $\tau$ the coordinate along $S^{1}$ and $Y_{r}=-I_{r}K$ are the left-invariant vector fields on $S^{3}$, where $I_{r}$ are the left-invariant complex structures on $S^{3}\times S^{1}$, see [23] for a detailed analysis. In more than four dimensions, $k>1$, a general structure theorem has recently been proved for the Lie algebra of the vector fields $Y_{r}$ in [53]. But this symmetry cannot automatically be lifted on $\mathscr{M}^{*}_{\mathrm{HE}}$, where the Hermitian-Einstein condition is taken with respect to one of the complex structures of the underlying HKT geometry, as the geometry of the $\mathscr{M}^{*}_{\mathrm{HE}}$ depends on the choice of complex structure and the vector fields $Y_{r}$ are not holomorphic with respect to $I_{s}$ for $r\not=s$.
@@
If in addition the HKT manifold $M^{4k}$ is compact, then as a consequence of Proposition 5.2, one has that $K\Phi=Y_{r}\Phi=0$. In four dimensions, $M^{4}$ is parallelisable because $K$ and $Y_{r}$ are also $^\widehat{\nabla}$-covariantly constant. Since $K$ commutes with $Y_{r}$ and the commutator is given by the components of $H$, $H$ can be written as $H=f\,Y^{\flat}_{1}\wedge Y^{\flat}_{2}\wedge Y^{\flat}_{3}$ for some function $f$. However, $H$ is invariant under the action of $K$ and $Y_{r}$, because $K$ is tri-holomorphic and Killing and $Y_{r}$ is holomorphic with respect to $I_{r}$ and Killing, which implies that $f$ is constant and $H$ is necessarily a closed 3-form. As $\Phi$ is invariant under the action of all four vectors fields $K$ and $Y_{r}$, $\Phi$ is constant. These manifolds for $H\not=0$ are identifications of $S^{3}\times S^{1}$ with a discrete group. Another way to state this result is that the unique set of hyper-complex structures on $S^{3}\times S^{1}$ [52] do not admit HKT geometries with $dH\not=0$ and non-constant dilaton $\Phi$. For a detailed description of the HKT structures on $S^{3}\times S^{1}$ see [23].